OpenStudy (anonymous):

A certain city block is in the form of parallelogram. Two of its sides measures 32ft and 41 ft. If the area of the land in the block is 656 ft^2, what is the length of its longer diagonal?

OpenStudy (zpupster):

parrellelogram Area A=b*h 

OpenStudy (zpupster):

we know A = 656 ft^2, b = 41 so we need to find h 656=41*h h=??

OpenStudy (anonymous):

h=16 Then?

OpenStudy (zpupster):

@klionmaster10

OpenStudy (zpupster):

pythorgorean theorem -- a^2 + b^2 = c^2 16^2 + x^2 = 32^2 32^2 - 16^2 =?? x=sqrt(32^2 - 16^2)

OpenStudy (anonymous):

Thanks!

OpenStudy (zpupster):

70.55 feet. this was your answer?? good job!!

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

how did you come up with 70.55 feet?

OpenStudy (anonymous):

already got h=16,